Find the point(s) where the tangent to the curve is horizontal for:
y=2(x-29)(x+1)

Answer by stanbon(75887)

Find the point(s) where the tangent to the curve is horizontal for:
y=2(x-29)(x+1)
---------------------
Take the derivative to get:
y' = 2[(x-29)*1 + (x+1)*1]
y' = 2[2x-28)
---------
Let y' = 0 and solve for x:
2[2x-28]= 0
x = 14
-------------------
The slope is zero when x = 14
------------------------------
f(14) = 2(14-29)(14+1) = 2(-15)(15) = -450
=================================
Ans: (14,-450)
======================
Cheers,
Stan H.